# Photoelectron Circular Dichroism of Aqueous-Phase Alanine

### v1

@@ 1 Introduction
In 2001, it was first experimentally demonstrated that one-photon photoionization of chiral molecules by circularly polarized synchrotron radiation may also be used to differentiate between enantiomers 4. This photoelectron circular dichroism (PECD) — predicted theoretically 25 years earlier by Ritchie 5, 6 — manifests as a forward-backward asymmetry in the measured photoelectron flux along an experimental axis defined by the direction of light propagation. Arising purely due to electric-dipole interactions, PECD is much more pronounced than ECD, exhibiting asymmetries on the order of 1% up to a few 10’sfor randomly oriented molecules 7. Alignment of molecules has been demonstrated to further increase PECD effect sizes 8, 9, 10. PECD appears to be a general feature in photoionization of chiral systems by CPL, having been observed in core-level and valence-band photoemission from terpenes 11, 12, 13, 14 as well as from molecular dimers and larger clusters 15, 16, amino acids 17, 18, 19, organometallic complexes 20, and nanoparticles 18. In addition to the traditional strengths of photoelectron spectroscopy, including chemical-state and site specificity, PECD is uniquely sensitive to small differences in molecular electronic structure and is capable of clearly distinguishing between different structural and conformational isomers of a chiral molecule 21, 22. PECD is a powerful method capable of much more than the resolution of enantiomers and the determination of enantiomeric excess. These very appealing analytical capabilities have driven the growing field of laser-based multi-photon PECD since 2012.23, 24, 25
@@
Despite these challenges, we have previously demonstrated the viability of liquid-phase PECD measurements for the chiral molecule fenchone as a neat liquid 30. We now report PECD in core-level photoionization of aqueous-phase alanine — the simplest chiral amino acid — at different solution pH values corresponding to the cationic, zwitterionic, and anionic form of the molecule. We summarize the results of our measurements, highlight the key differences in measuring PECD in the aqueous and gas phases, and discuss the potential influence of water solvation on PECD. We outline new opportunities for aqueous-phase PECD going forward, including the relatively straightforward application of PECD to study cationic or anionic chiral molecules; a topic that has only recently begun to be explored 35, 36, 37. More generally, this study constitutes the first core-level PECD investigation of chiral amino acids in any context. Liquid-phase PECD studies of the elementary constituents of life are highly relevant to in vivo biological conditions. We present a first step toward a bottom-up approach to understanding biomolecular complexity through the preparation of simple aqueous solutions, enabling the study of fragile thermolabile chiral species such as amino acids, peptides, sugars, and nucleic acids, whose vaporization into the gas phase is extremely challenging.
@@ 1.1 Experimental Considerations
In this work, we focus on core-level C 1s measurements of aqueous-phase alanine. Alanine has three chemically distinct carbon atoms — a carboxylic acid group, a chiral central carbon adjacent to an amine group, and a methyl group, which we denote C1, C2, and C3, respectively — resulting in a solution-phase C 1s PE spectrum with three distinct features that overlap to varying degrees depending on the solution pH / alanine protonation state (Fig. 1). An investigation of the precisely energy-referenced C 1s spectrum of alanine is not the focus of this work, and we have aligned all spectra discussed below to the C3 peak position (corresponding the alanine’s methyl group, see gray dashed line in Fig. 1) in order to better illustrate the kinetic-energy (KE) shifts occurring for C1 and C2 upon protonation or deprotonation. As we did not employ the energy-referencing scheme outlined in our earlier work 38, 39, 40, the KE axes presented in this work represent as-measured values. These values may be subject to inaccuracies due to spectrometer offsets, the presence of streaming potentials and other stray fields, and changes in the bulk electronic structure of water upon the addition of large amounts of NaOH and HCl needed for pH adjustment 41, 42, 43.
To track chemical shifts upon protonation / deprotonation of alanine’s functional groups, we first measured PE spectra from aqueous-phase alanine for solution pH values of 1, 7, and 13 using 335 eV photons, thereby generating photoelectrons with approximately 40–45 eV KE. These pH values were chosen based on the pKa values of alanine’s functional groups (2.3 and 9.9 for alanine’s carboxylic acid and amine groups in water, respectively) and ensured that, for a given solution pH, >95% (>99% for pH 7 and 13) of the alanine molecules in solution adopted a single charge state. At these rather high KEs the primary PE features are well separated from the high-intensity background region mentioned above, which becomes significant below approximately 20 eV KE.
@@
Under strongly acidic conditions, alanine is predominantly in a cationic state ($Ala^{+}$), with the amine functional group protonated (Fig. 1, top). In this state, despite some overlap between C2 and C3, the three carbon PE features are easily resolvable owing to pronounced chemical shifts arising due to the different chemical environments of the carbon sites. The lowest KE feature corresponds to photoelectrons originating from the relatively electronegative COOH group (C1). The central PE feature corresponds to the chiral center adjacent to the amine group (C2), and the final feature is attributable to the methyl group (C3). These assignments are consistent with previous solid-state and gas-phase work 44, 45. As can be seen by comparing the pH 1 and pH 7 data in Figure 1, deprotonation of the carboxylic acid for zwitterionic Alazw is accompanied by a noticeable chemical shift of the C1 peak (relative to the C3 peak) of approximately 0.71 eV toward higher KE (lower binding energy), reflecting the increasing electron density at that site upon deprotonation. In contrast, C2 does not experience much of a chemical shift, suggesting that the central carbon is not particularly sensitive to the changing electron density on the neighboring functional group. This may be a consequence of effective charge screening by nearby water molecules.46 At this pH, all three primary PE features are still easily resolvable. At pH 13, alanine is found in the anionic form, $Ala^{-}$, with the carboxylic acid deprotonated and the amine group neutral. In this case, we observe a chemical shift of 0.44 eV toward lower KE for the C2 peak, corresponding to an increase in electron density at that group compared to the pH 7 case. This chemical shift results in substantial overlap of the C2 and C3 peaks. We note that the C1 peak remains clearly isolated from the C2 and C3 features for all charge states, thereby simplifying the determination of PECD for this carbon site.
@@
Figure 1: Representative as-measured C 1s photoelectron spectra corresponding to aqueous-phase alanine’s three chemically distinct carbons measured with $h\nu$ = 335 eV photons under acidic (top), neutral (middle), and basic (bottom) aqueous conditions. Measurements were conducted with 1 M aqueous solutions of L-alanine. The dashed black curves are fits to the data following the subtraction of background signals (dashed purple lines) using Gaussian profiles, with the widths of the overlapping C2 and C3 peaks constrained to differ by a maximum of 10%. All spectra were shifted to the as-measured kinetic energy of the C3 peak, corresponding to the methyl group (dashed gray vertical line). Molecular structures of the dominant protonation state of alanine at the various pH values investigated are shown as insets.
@@
For single-photon photoionization of a randomly oriented sample 47, the normalized photoelectron angular distribution (PAD) within the electric-dipole approximation is given by:
@@
Here, $I^{p}(\theta)$ is the photoionization intensity at an angle $\theta$ measured with respect to the photon-propagation direction for a given light polarization ($p$ = ±1 for circularly polarized light and 0 for linearly polarized light, respectively). $P_{n}$ represents the Legendre polynomial of the nth order. The two coefficients $b^{p}_{1}$ and $b^{p}_{2}$ completely encompass the target-specific contribution to the PAD under a fixed set of experimental conditions. For achiral molecules, $b^{p}_{1}=0$ and thus the PAD is described by $b^{p}_{2}$, more commonly denoted $\beta$ for experiments involving linearly polarized light (i.e. $b^{0}_{2}=\beta$, while $b^{\pm 1}_{2}=-\beta/2$) 48. For chiral molecules, $b^{\pm 1}_{1}$ may be non-zero and $b^{+1}_{1}=-b^{-1}_{1}$, changing sign upon inversion of the handedness of CPL (or of the enantiomer). An asymmetry factor, $G$, may be defined as the simple difference between left- and right-handed PE intensities divided by their average. Using Eq. 1 and noting that $P_{1}(x)=x$, we then have:
@@
$b^{+1}_{1}=\pm\frac{[I^{+1}-I^{-1}]}{[I^{+1}+I^{-1}]}\cdot\frac{1}{cos(54.74^{\circ})},$ (3)
@@
As $b^{+1}_{1}$ simply changes sign upon exchange of enantiomer, in principle only a single pair of measurements for a single enantiomer with left- and right-handed CPL is necessary for its determination at any given energy using Eq. 2. The normalized intensities $I^{p}(\theta)$ are obtained from the experimental photoelectron intensities at a specified electron energy and detection angle $\theta$, but require correction accounting for possible variation in photon flux and other sources of medium term experimental drift. Despite our best efforts to ensure that the fluxes of left- and right-handed CPL were equal — monitored via a photodiode — small persistent differences in light intensity of 1–3% remained and so data normalization was essential. In gas-phase PECD studies, $4\pi$ imaging spectrometers enable collection of the total ionization yield, and thus enable direct normalization.12 In the present experimental arrangement the total ionization yield cannot be monitored and an alternative intensity-correction procedure utilizing the baseline signal was adopted. It is assumed that the background of scattered electrons is achiral and essentially independent of light polarization state. We therefore scaled all left- and right-handed CPL data (hereafter denoted as $+$ and -- spectra) to ensure equal intensity of the spectra at the low- and high-KE edges, where only an achiral background is present.
@@
It is also vital that the baseline is accurately modeled such that the background beneath the peaks of interest can be reliably removed from the intensities $I$ used to determine $b^{+1}_{1}$. It can be seen from Eq. 2 that while any residual background present in the suitably scaled $I^{+}$ and $I^{-}$ will self-cancel in the numerator, this does not apply to the denominator, $[I^{+}+I^{-}]$, whose magnitude would hence be overestimated, and lead to an erroneous underestimation of the derived $|b^{+1}_{1}|$ (see Fig. 3). This challenge is exacerbated for experiments with dilute chiral molecules where the signal-to-background ratio is much reduced (see Fig. 2, top). Details regarding the background-subtraction method employed may be found in the supporting information.
@@
At the near magic-angle geometry of $\theta=50^{\circ}$ (in the backward direction) adopted for these experiments, the polynomial $P_{2}(\cos\theta)\neq 0$ and some residual influence of the $b^{\pm 1}_{2}$ parameter is unavoidably included in the measurement (see Eq. 2). However, there are currently no published values of $b^{\pm 1}_{2}$ for core-level photoionization of alanine. To obtain an estimate of the significance of the remnant $b^{\pm 1}_{2}$ contribution, we calculated the value of $b^{0}_{2}$ ($\beta$) for low-energy (KE = 0 to 25 eV) photoelectrons generated upon C 1s photoionization of gas-phase $Ala^{0}$ (details may be found in the supporting information, see Fig. LABEL:SI-fig:S2). We found that for photoionization of C1, the absolute value of $\beta$ is less than 0.8 within this entire KE range, decreasing to less than 0.5 for photoelectrons with KE < 15 eV. Although aqueous-phase alanine will undoubtedly yield different values of $\beta$ due to differences in conformation and the charge state of its functional groups, this calculation nevertheless provides a rough estimate. As a comparison, the measured value of $\beta$ for O 1s photoionization of gas-phase water at 25 eV KE is approximately 1.6, whereas for liquid water at the same energy it is 0.6 34. For both environments, $\beta$ trends strongly toward 0 for lower photoelectron kinetic energies. This decrease in $\beta$ as measured from liquid versus gaseous water was attributed primarily to highly efficient elastic or quasi-elastic scattering of photoelectrons by water molecules in the liquid. As water remains by far the likeliest scatterer of photoelectrons for the 1 M aqueous solutions of alanine studied in this work, it seems reasonable to expect a similar degree of PAD isotropization for $\beta$ as measured in C 1s photoionization of these solutions. Conservatively assuming a $\beta$ value of 0.5 as an upper limit for aqueous-phase alanine and using Eq. 2 while recalling that $b^{\pm 1}_{2}=-\beta/2$ and simply averaging over the acceptance angle of the analyzer (±$\approx\pm 15^{\circ}$) 29, the effect would be to scale our measured values of $b^{+1}_{1}$ by a factor of 0.97. Such a change does not alter the results of this study, and we do not consider it further in subsequent discussion.

### v2

@@ 1 Introduction
In 2001, it was first experimentally demonstrated that one-photon photoionization of chiral molecules by circularly polarized synchrotron radiation may also be used to differentiate between enantiomers 4. This photoelectron circular dichroism (PECD) — predicted theoretically 25 years earlier by Ritchie 5, 6 — manifests as a forward-backward asymmetry in the measured photoelectron flux along an experimental axis defined by the direction of light propagation. Arising purely due to electric-dipole interactions, PECD is much more pronounced than ECD, exhibiting asymmetries on the order of 1% up to a few 10’s% for randomly oriented molecules 7. Alignment of molecules has been demonstrated to further increase PECD effect sizes 8, 9, 10. PECD appears to be a general feature in photoionization of chiral systems by CPL, having been observed in core-level and valence-band photoemission from terpenes 11, 12, 13, 14 as well as from molecular dimers and larger clusters 15, 16, amino acids 17, 18, 19, organometallic complexes 20, and nanoparticles 18. In addition to the traditional strengths of photoelectron spectroscopy, including chemical-state and site specificity, PECD is uniquely sensitive to small differences in molecular electronic structure and is capable of clearly distinguishing between different structural and conformational isomers of a chiral molecule 21, 22. PECD is a powerful method capable of much more than the resolution of enantiomers and the determination of enantiomeric excess. These very appealing analytical capabilities have driven the growing field of laser-based multi-photon PECD since 2012.23, 24, 25
@@
Despite these challenges, we have previously demonstrated the viability of liquid-phase PECD measurements for the chiral molecule fenchone as a neat liquid 30. We now report PECD in core-level photoionization of aqueous-phase alanine — the simplest chiral amino acid — at different solution pH values corresponding to the cationic, zwitterionic, and anionic form of the molecule. We summarize the results of our measurements, highlight the key differences in measuring PECD in the aqueous and gas phases, and discuss the potential influence of water solvation on PECD. We outline new opportunities for aqueous-phase PECD going forward, including the relatively straightforward application of PECD to study cationic or anionic chiral molecules; a topic that has only recently begun to be explored 35, 36, 37. More generally, this study constitutes the first core-level PECD investigation of chiral amino acids in any context. Liquid-phase PECD studies of the elementary constituents of life are highly relevant to in vivo biological conditions. We present a first step toward a bottom-up approach to understanding biomolecular complexity through the preparation of simple aqueous solutions, enabling the study of fragile thermolabile chiral species such as amino acids, peptides, sugars, and nucleic acids, whose vaporization into the gas phase with sufficient density to enable core-level PES is extremely challenging.
@@ 1.1 Experimental Considerations
In this work, we focus on core-level C 1s measurements of aqueous-phase alanine. Alanine has three chemically distinct carbon atoms — a carboxylic acid group, a chiral central carbon adjacent to an amine group, and a methyl group, which we denote C1, C2, and C3, respectively — resulting in a solution-phase C 1s PE spectrum with three distinct features that overlap to varying degrees depending on the solution pH / alanine protonation state (Fig. 1).
To track chemical shifts upon protonation / deprotonation of alanine’s functional groups, we measured C 1s PE spectra for aqueous solutions of alanine with solution pH values of 1, 6, and 13 using 480.46 eV photons. These pH values were chosen based on the $p\textit{K}_{a}$ values of alanine’s functional groups (2.3 and 9.9 for alanine’s carboxylic acid and amine groups in water, respectively) and ensured that, for a given solution pH, >95% (>99% for pH 6 and 13) of the alanine molecules in solution adopted a single charge state. For these measurements, we employed the energy-referencing scheme that we have developed over the past years.38, 39, 40 Briefly, this method enabled us to correct for the influence of electrokinetic charging and other stray fields by applying a suitable bias voltage to the liquid jet prior to measurement, thereby increasing the kinetic energy of all photoelectrons commensurately with the bias applied and revealing the secondary electron cutoff of the spectrum, which represents a true zero kinetic-energy reference for condensed-phase experiments. We note that it was not practical to employ these measures for the low-kinetic-energy PECD measurements discussed later, and thus the axes displayed in subsequent figures represent as-measured values, which may be subject to inaccuracies due to spectrometer offsets and the presence of streaming potentials and other stray fields 41, 42.
@@
Under strongly acidic conditions, alanine is predominantly in a cationic state ($Ala^{+}$), with the amine functional group protonated (Fig. 1, top). In this state, despite some overlap between C2 and C3, the three carbon PE features are easily resolvable owing to pronounced chemical shifts arising due to the different chemical environments of the carbon sites. The highest binding energy (BE) feature corresponds to photoelectrons originating from the relatively electronegative COOH group (C1). The central PE feature corresponds to the chiral center adjacent to the amine group (C2), and the final feature is attributable to the methyl group (C3). These assignments are consistent with previous solid-state and gas-phase work 43, 44. As can be seen by comparing the pH 1 and pH 6 data in Fig. 1, deprotonation of the carboxylic acid for zwitterionic Alazw is accompanied by a noticeable chemical shift of the C1 peak of approximately 1.12 eV toward lower BE, reflecting the increasing electron density at that site upon deprotonation. C2 and C3 also exhibit shifts to lower BE, albeit of smaller magnitudes, indicating their relatively lower sensitivity to the changing electron density on the neighboring functional group. This is likely a consequence of effective charge screening by nearby water molecules.45 At this pH, all three primary PE features are still readily resolvable. At pH 13, alanine is found in the anionic form, $Ala^{-}$, with the carboxylic acid deprotonated and the amine group neutral. In this case, we observe a chemical shift of 0.77 eV toward lower BE for the C2 peak, corresponding to an increase in electron density at that group compared to the pH 6 case. This chemical shift results in substantial overlap of the C2 and C3 peaks. We note that the C1 peak remains clearly isolated from the C2 and C3 features for all charge states, thereby simplifying the determination of PECD for this carbon site. The binding energies for C1, C2, and C3 for solution pH 1, 6, and 13 are summarized in Table 1
Given the relatively high kinetic energy ($\approx 190~eV$) of the photoelectrons measured, these C 1s measurements were not particularly surface sensitive. Nevertheless, by comparing the relative peak area of C1, C2, and C3, we can draw initial inferences about the orientation of surface-localized molecules. Normalizing to the area of C1, we find that the C1:C2:C3 ratios are 1:1.09:1.13, 1:1.12:1.09, and 1:1.23:1.07 for Ala+, Alazw, and Ala-, respectively. Assuming that the C 1s photoionization cross section differs only negligibly for C1, C2, and C3, a completely random molecular orientation distribution would result in a peak-area ratio of 1:1:1. A deviation from this ratio suggests partial molecular orientation for molecules near the liquid-vapor interface, with lower peak areas indicating that the photoionized groups are located deeper within the solution bulk 46. Our results suggest that C1 appears to reside further in the solution bulk than C2 or C3 at all solution pH values probed. The disparity in the C1:C2 ratio is smallest for Ala+ and most significant for Ala-, where the charged carboxylic group is likely to be fully solvated. However, due to its overall hydrophilicty, alanine does not exhibit strong surface propensity 47, and oriented molecules near the liquid-vapor interface are likely to make up only a fraction of the total molecules probed. Nevertheless, these potential effects are important to keep in mind, particularly for PECD measurements, which inevitably involve low-kinetic-energy electrons and are thus particularly surface sensitive.
@@
Figure 1: Representative as-measured C 1s photoelectron spectra corresponding to aqueous-phase alanine’s three chemically distinct carbons measured with $h\nu$ = 480.46 eV photons under acidic (top), neutral (middle), and basic (bottom) aqueous conditions. Measurements were conducted with 1 M aqueous solutions of L-alanine. The dashed black curves are fits to the data following the subtraction of background signals (dashed purple lines) using exponentially modified Gaussian profiles 48. The asymmetry parameter was constrained to be between 0.2 and 0.3, and these values were kept constant in fits across different pH conditions. The ratios of peak widths were also constrained to differ by no more than 5% across the three fits. Molecular structures of the dominant protonation state of alanine at the various pH values investigated are shown as insets.
Table 1: C 1s binding energies for each of alanine’s carbon groups as a function of solution pH.
Solution pH BE C1 [eV] BE C2 [eV] BE C3 [eV]
1 294.64 292.12 290.53
6 293.52 291.56 290.19
13 293.25 290.79 289.81
@@
For single-photon photoionization of a randomly oriented sample 49, the normalized photoelectron angular distribution (PAD) within the electric-dipole approximation is given by:
@@
Here, $I^{p}(\theta)$ is the photoionization intensity at an angle $\theta$ measured with respect to the photon-propagation direction for a given light polarization ($p$ = ś1 for circularly polarized light and 0 for linearly polarized light, respectively). $P_{n}$ represents the Legendre polynomial of the nth order. The two coefficients $b^{p}_{1}$ and $b^{p}_{2}$ completely encompass the target-specific contribution to the PAD under a fixed set of experimental conditions. For achiral molecules, $b^{p}_{1}=0$ and thus the PAD is described by $b^{p}_{2}$, more commonly denoted $\beta$ for experiments involving linearly polarized light (i.e. $b^{0}_{2}=\beta$, while $b^{\pm 1}_{2}=-\beta/2$) 50. For chiral molecules, $b^{\pm 1}_{1}$ may be non-zero and $b^{+1}_{1}=-b^{-1}_{1}$, changing sign upon inversion of the handedness of CPL (or of the enantiomer). An asymmetry factor, $G$, may be defined as the simple difference between left- and right-handed PE intensities divided by their average. Using Eq. 1 and noting that $P_{1}(x)=x$, we then have:
@@
$b^{+1}_{1}=\pm\frac{[I^{+1}-I^{-1}]}{[I^{+1}+I^{-1}]}\times\frac{1}{cos(54.74^{\circ})},$ (3)
@@
As $b^{+1}_{1}$ simply changes sign upon exchange of enantiomer, in principle only a single pair of measurements for a single enantiomer with left- and right-handed CPL is necessary for its determination at any given energy using Eq. 2. The normalized intensities $I^{p}(\theta)$ are obtained from the experimental photoelectron intensities at a specified electron energy and detection angle $\theta$, but require correction accounting for possible variation in photon flux and other sources of medium term experimental drift. Despite our best efforts to ensure that the fluxes of left- and right-handed CPL were equal — monitored via a photodiode — small persistent differences in light intensity of 1–3% remained and so data normalization was essential. In gas-phase PECD studies, $4\pi$ imaging spectrometers enable collection of the total ionization yield, and thus enable direct normalization.12 In the present experimental arrangement the total ionization yield cannot be monitored and an alternative intensity-correction procedure utilizing the baseline signal was adopted. It is assumed that the background of scattered electrons is achiral and essentially independent of light polarization state. We therefore scaled all left- and right-handed CPL data (hereafter denoted as ++ and - spectra) to ensure equal intensity of the spectra at the low- and high-KE edges, where only an achiral background is present.
@@
It is also vital that the baseline is accurately modeled such that the background beneath the peaks of interest can be reliably removed from the intensities $I$ used to determine $b^{+1}_{1}$. It can be seen from Eq. 2 that while any residual background present in the suitably scaled $I^{+}$ and $I^{-}$ will self-cancel in the numerator, this does not apply to the denominator, $I^{+}+I^{-}]$, whose magnitude would hence be overestimated, and lead to an erroneous underestimation of the derived $|b^{+1}_{1}|$. This challenge is exacerbated for experiments with dilute condensed-phase chiral molecules where the signal-to-background ratio is much reduced (see Fig. 2, top). Details regarding the background-subtraction method employed may be found in the supplementary information.
@@
At the near magic-angle geometry of $\theta=50^{\circ}$ (in the backward direction) adopted for these experiments, the polynomial $P_{2}(\cos\theta)\neq 0$ and some residual influence of the $b^{\pm 1}_{2}$ parameter is unavoidably included in the measurement (see Eq. 2). However, there are currently no published values of $b^{\pm 1}_{2}$ for core-level photoionization of alanine. To obtain an estimate of the significance of the remnant $b^{\pm 1}_{2}$ contribution, we calculated the value of $b^{0}_{2}$ ($\beta$) for low-energy (KE = 0 to 25 eV) photoelectrons generated upon C 1s photoionization of gas-phase $Ala^{0}$ (details may be found in the supplementary information, see Fig. LABEL:SI-fig:S2). We found that for photoionization of C1, the absolute value of $\beta$ is less than 0.8 within this entire KE range, decreasing to less than 0.5 for photoelectrons with KE < 15 eV. Although aqueous-phase alanine will undoubtedly yield different values of $\beta$ due to differences in conformation and the charge state of its functional groups, this calculation nevertheless provides a rough estimate. As a comparison, the measured value of $\beta$ for O 1s photoionization of gas-phase water at 25 eV KE is approximately 1.6, whereas for liquid water at the same energy it is 0.6 34. For both environments, $\beta$ trends strongly toward 0 for lower photoelectron kinetic energies. This decrease in $\beta$ as measured from liquid versus gaseous water was attributed primarily to highly efficient elastic or quasi-elastic scattering of photoelectrons by water molecules in the liquid. As water remains by far the likeliest scatterer of photoelectrons for the 1 M aqueous solutions of alanine studied in this work, it seems reasonable to expect a similar degree of PAD isotropization for $\beta$ as measured in C 1s photoionization of these solutions. Conservatively assuming a $\beta$ value of 0.5 as an upper limit for aqueous-phase alanine and using Eq. 2 while recalling that $b^{\pm 1}_{2}=-\beta/2$ and simply averaging over the acceptance angle of the analyzer ($\approx\pm 15^{\circ}$) 29, the effect would be to scale our measured values of $b^{+1}_{1}$ by a factor of 0.97. Such a change does not alter the results of this study, and we do not consider it further in subsequent discussion.